OpenStudy (anonymous):

find a so that the point (-1,2) is on the graph of f(x)=ax^2+5

OpenStudy (anonymous):

f(x) be y. Sub (-1,2) to equation. 2=a(-1)^2+5 2=a+5 a=2-5 a=-3

OpenStudy (anonymous):

thank you so much!! I feel really dumb right about now!

OpenStudy (anonymous):

Welcome! It's okay! Goodluck!:)